Consider a power plant is operating With coal combustion by producing temperature sum of 20.75 0F, 20.4K, 289.ec,  30.450F, 500.850 F and 200.15K, while converting the heat content of power plant in useful work with an average of 200.550F, 200K and 300.450C„ then find the heat dissipated in the vicinity and also find the % efficiency of the power plant


1
Expert's answer
2021-01-29T16:57:08-0500

"T_1=266.9+20.4+562.15+272.29+533.62+200.15=1855.5~K."

"T_2=366.79+200+573.6=1140.4~K."

"\eta=1-\frac{T_2}{T_1}=38.54~" %.